OpenStudy (anonymous):

Find each union or intersection. Let A = {1,3,4}, B = {x|x is an even whole number less than 9}, C = {2,5,7,10} and D = {x|x is an odd whole number less than 10}. 1. C U D 2. A (upside down U) D

OpenStudy (mathmath333):

so \(\large\tt \begin{align} \color{black}{ A=\{ 1,3,4 \}\hspace{.33em}\\~\\ B=\{x|\text{x is an even whole number less than 9}\}\hspace{.33em}\\~\\ B=\{0,2,4,6,8\}\hspace{.33em}\\~\\ C = \{2,5,7,10\}\hspace{.33em}\\~\\ D = \{x|\text{x is an odd whole number less than 10}\}\hspace{.33em}\\~\\ D = \{1,3,5,7,9\}\hspace{.33em}\\~\\ }\end{align}\) see if it makes sense

OpenStudy (anonymous):

i gothca

OpenStudy (mathmath333):

so what is \(\large\tt \begin{align} \color{black}{ C\cup D\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

i don't know what the sign means exactly

OpenStudy (mathmath333):

it means all values of \(C\) and \(D\)

OpenStudy (anonymous):

{1,2,3,5,7,9,10} ?

OpenStudy (mathmath333):

\(\Huge \checkmark\)

OpenStudy (anonymous):

Thanks ! What about #2 ?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ A\cap D \hspace{.33em}\\~\\ }\end{align}\) means only values which are in both \(A\) and \(D\)

OpenStudy (anonymous):

{1,3} ?

OpenStudy (mathmath333):

\(\Huge \checkmark\)

OpenStudy (anonymous):

THANK YOU !!!